OpenStudy (anonymous):

what is the solution of k+4/2=k+9/3

OpenStudy (anonymous):

no solution empty set {}

OpenStudy (gw2011):

If you mean: (k+4)/2=(k+9)/3, then the solution is: Multiply both sides by 6 in order to get rid of the denominator: 3(k+4)=2(k+9) 3k+12=2k+18 k=6 However, if you mean: k+(4/2)=k+(9/3), then there is no solution.

OpenStudy (anonymous):

hmm i didn't think of that good call you're probably right